find the equation of the line though (1,4) and slope 1/4
Answer by troyapplehelen(46)

the formula of a line is y=mx+b, and here m is 1/4
so, y=(1/4)x+b and this line passes the point (1,4)
so, 4=(1/4)times(1)+b, which is 4=(1/4)+b. so then b is 15/4.
the answer, y=(1/4)x+15/4.